# Adaptive and Reliable Clinical Reasoning

### v2

@@ Abstract
In high-stakes domains like clinical reasoning, AI assistants powered by large language models (LLMs) are yet to be reliable and safe. We identify a key obstacle towards reliability: existing LLMs are trained to answer any question, even with incomplete context in the prompt or insufficient parametric knowledge. We propose to change this paradigm to develop more careful LLMs that ask follow-up questions to gather necessary and sufficient information and respond reliably. We introduce MediQ, a framework to simulate realistic clinical interactions, which incorporates a Patient System and an adaptive Expert System. The Patient may provide incomplete information in the beginning; the Expert refrains from making diagnostic decisions when unconfident, and instead elicits missing details from the Patient via follow-up questions. To evaluate MediQ, we convert MedQA and Craft-MD—medical benchmarks for diagnostic question answering—into an interactive setup. We develop a reliable Patient system and prototype several Expert systems, first showing that directly prompting state-of-the-art LLMs to ask questions degrades the quality of clinical reasoning, indicating that adapting LLMs to interactive information-seeking settings is nontrivial. We then augment the Expert with a novel abstention module to better estimate model confidence and decide whether to ask more questions, thereby improving diagnostic accuracy by 22.3%; however, performance still lags compared to an (unrealistic in practice) upper bound when full information is given upfront. Further analyses reveal that interactive performance can be improved by filtering irrelevant contexts and reformatting conversations. Overall, our paper introduces a novel problem towards LLM reliability, a novel MediQ framework, and highlights important future directions to extend the information-seeking abilities of LLM assistants in critical domains.
@@ 1 Introduction
General-purpose LLMs are designed to serve a broad audience by following instructions and providing the most likely and general answers (Brown et al., 2020; Ouyang et al., 2022). However, in high-stakes decision making scenarios such as medical conversations, LLM assistants can be harmful if they provide general responses instead of gathering missing information to make informed decisions. As shown in Figure 1, standard medical question-answering (QA) tasks are formulated in a single-turn setup where all necessary information is provided upfront, and the model is not expected to interact with users (Jin et al., 2021; Pal et al., 2022; Jin et al., 2019; Hendrycks et al., 2020). This QA paradigm deviates from realistic healthcare environments where patients often provide incomplete information at intake, and clinicians arrive at the diagnosis via an investigative process (Trimble & Hamilton, 2016; Bornstein & Emler, 2001; Masic, 2022).
@@
This gap between existing benchmarks and reality calls for a paradigm shift to designing systems adept at navigating high-stakes interactive scenarios. Focusing on clinical interactions as a primary use case, we introduce MediQ, an interactive framework for medical evaluation with dynamic information-seeking questions, to address the static limitations of current medical QA benchmarks (Figure 2). Unlike conventional systems, which assume that all necessary information is readily available, MediQ acknowledges the inherent uncertainty in medical consultations where a typical patient does not have the expertise to distill all necessary and relevant information they need to provide. To achieve this, MediQ comprises two components: a Patient system that simulates a patient and responds to follow-up questions, and an Expert system that serves as a doctor’s assistant and asks questions to the patient before making a medical decision. In this interactive clinical reasoning task, a successful information-seeking Expert should decide, at each turn, whether it has enough information to provide a confident answer; if not, it should ask a follow-up question.
We convert two medical QA datasets, MedQA (Jin et al., 2021) and Craft-MD (Johri et al., 2023, 2024), into an interactive setup by parsing the patient records to only provide partial information in the beginning, allowing the Patient to withhold details until requested by the Expert. We first develop and validate a GPT-3.5-based Patient system that can accurately answer Expert inquiries by retrieving the correct facts from the patient record. We then benchmark Expert systems based on state-of-the-art (SOTA) LLMs, including Llama-3 (Touvron et al., 2023), GPT-3.5 (Brown et al., 2020) and GPT-4 (OpenAI et al., 2024), to evaluate their proactive information seeking ability. It is striking that their average accuracy drops by a relative 11.3% compared to starting with the same limited information and asking no questions, showing that adapting LLMs to interactive information-seeking settings is nontrivial. A key challenge is deciding when to ask a follow-up question instead of directly providing an answer. Augmenting it with confidence estimation strategies with explicit reasoning and self-consistency, we improve GPT-3.5-based Expert system performance across datasets by a relative 22.3%; however, it still lags by 10.3% compared to an upper bound when directly given full information.
Our results show that while SOTA LLMs perform relatively well when given complete information, they struggle to proactively seek information in a realistic interactive setting with incomplete initial information. By providing a modular, interactive benchmark, we hope to facilitate the development of reliable LLM assistants for realistic clinical reasoning scenarios. Our main contributions are:
@@
    We introduce the MediQ framework, simulating realistic medical consultations with modular components to improve clinical reasoning in LLMs. The framework consists of a Patient system and an adaptive Expert system that interacts with the Patient to perform clinical reasoning.
@@
    We develop a reliable Patient system that interacts with any Expert system, evaluating the Expert’s medical information seeking and clinical reasoning abilities.
@@
    We show that SOTA LLMs such as Llama-3-Instruct and GPT-4 struggle at proactive information seeking, revealing a significant gap in this area.
@@
    We develop MediQ-Expert, our best Expert system with novel abstaining capabilities to reduce unconfident answers, to partially close the gap between the realistic incomplete information setup and the existing full information setup. We make our code and data publicly available. [^1] https://github.com/stellali7/MediQ
@@ 2 MediQ: Dynamic Medical Consultation Framework Overview
Figure 2: The MediQ Framework. MediQ operationalizes a realistic dynamic medical consultation process. It consists of a Patient system and an Expert system to evaluate medical info-seeking.
@@ 2.2 The Expert System
Expert Task    The Expert system simulates the medical decision-making process of experienced clinicians, who seek additional patient information and iteratively update their differential diagnosis. The Expert system is first presented with a medical question and limited patient information. As each turn, it assesses whether the provided information is sufficient to answer the question. If the Expert system is unconfident, it can elicit evidence with a follow-up information-seeking question; otherwise, the Expert system deems the acquired information sufficient and provides a final answer. The performance of the Expert system is evaluated on the efficiency of the conversation (number of follow-up questions) and the accuracy of the final diagnosis.
@@ 5.2 How does abstention improve Expert system performance?
Figure 8: Effect of the abstention module on Expert system performance (efficiency & accuracy). (a) Accuracy over # questions with no abstain response in question generator context (NC). Each line is an abstain method with increasing confidence thresholds. Accuracy increases with #Qs as threshold increases. (b) Expert confidence throughout conversation with and without rationale generation (RG), averaged over self-consistency levels 1, 3, and 5. RG leads to more conservation confidence estimates. (c) Accuracy over # questions comparing having abstention response in-context and not in-context (NC). In-context abstention improves accuracy, amplified with rationale generation (RG) and self-consistency (SC).
@@ 6 Related Work
Medical Question Answering Systems   Advancements in medical question answering (QA) systems have progressed from rule-based systems to LLM-powered agents. Notable medical QA benchmarks include MultiMedQA (Singhal et al., 2023a), which contains both multiple-choice and open-ended questions collected from various sources. To customize or adapt a general-purpose LLM to the medical domain, prior work often finetune the model on medical knowledge data such as PubMed (Bolton et al., 2022; Yasunaga et al., 2022; Wu et al., 2023a; Singhal et al., 2023a, b), or more recently on conversational medical datasets (Yunxiang et al., 2023; Han et al., 2023). Despite their proficiency in direct answer retrieval, the proactive information-seeking capability is not something these models are inherently designed to do. Our proposed methodological framework, MediQ, is designed to work as an overlay to these domain-specific models, enhancing them with the capability to actively seek additional information in a structured and clinically relevant manner.
Interactive Models and Agents   Interactive conversational models extend beyond the standard QA framework by engaging in a dialogue such as customer support and negotiation (Singh & Beniwal, 2022; Chakrabarti & Luger, 2015; He et al., 2018; Abdelnabi et al., 2023), where iterative information gathering is crucial. (Li et al., 2023) and (Andukuri et al., 2024) attempt to use LLMs to elicit more information-rich human preference examples in everyday tasks. However, the application of these models in the medical domain remains limited (Li et al., 2021). (Wu et al., 2023b) attempts to evaluate general-purpose LLMs and chain-of-thought reasoning on DDXPlus (Fansi Tchango et al., 2022), a rule-based synethtic patient dataset. (Hu et al., 2024) navigates the information-seeking scenario by developing a reward model guided by uncertainty and includes medical diagnosis as one of the tasks. (Johri et al., 2023) observed a similar phenomenon where LLM-doctors cannot elicit complete patient information, but do not focus on improving the information-seeking ability. The system proposed by (Tu et al., 2024) performs a multitude of medical tasks but does not explore the crucial problem of abstention. Lastly, multi-agent and human-AI collaboration frameworks have shown impressive interactive performance (Zhou et al., 2024; Wu et al., 2024, 2023c; Deng et al., 2024; Lin et al., 2024), and can greatly benefit from our novel interactive abstention methods to seek additional information. Our work fills this gap via providing a benchmark, a dataset, and a framework to comprehensively studying information-seeking abilities in clinical decision-making, and most importantly, opens the door for future endeavors in this direction.
@@ 7 Conclusion
In this paper, we identify a significant gap in current LLMs’ capability to ask clarifying questions and proactively seek information in settings where personalization, precision, and reliability are critical. We introduced MediQ, a framework that simulates realistic clinical interactions where the expert needs to seek more information. We showed that LLMs like Llama-3 and GPT-3.5 struggle to gather necessary information for accurate medical decisions. To address this problem, we presented novel Expert LLMs augmented with confidence judgment and medical expertise, substantially improving its clinical reasoning abilities. MediQ operationalizes interactive and explicit clinical reasoning processes, with added interpretability in the reasoning flow of language models and decision making. We encourage future research to refine the MediQ framework, expand medical knowledge integration, and customize the interactions to better serve the healthcare community.
@@
One limitation is the scarcity of datasets that contain detailed patient information sufficient for a medical diagnosis which, to the best of our knowledge, was only met by MedQA and Craft-MD. The majority of available medical datasets are designed to test models’ medical domain knowledge. Second, the Patient system in our benchmark relies on a paid API; future work should establish an open-source Patient. Lastly, our evaluation framework, while designed to be realistic, is still limited to the multiple-choice format. However, the flexibility of MediQ allows easy extensions into open-ended settings with appropriate datasets and well-defined conversation-level metric. Future work can focus on collecting a rich dataset in open-ended medical consultations and expanding the MediQ framework.
@@
If a similar framework is used in more realistic setups, users and clinicians should be trained to prevent the over-reliance on technology that is liable to make mistakes, and to understand its privacy and fairness risks.
@@ B.1 Initial Assessment
Given the information from above, your task is to choose one of four options that best answers the inquiry.

### v3

# MediQ: Question-Asking LLMs and a Benchmark for Reliable Interactive Clinical Reasoning

@@ Abstract
Users typically engage with LLMs interactively, yet most existing benchmarks evaluate them in a static, single-turn format, posing reliability concerns in interactive scenarios. We identify a key obstacle towards reliability: LLMs are trained to answer any question, even with incomplete context or insufficient knowledge. In this paper, we propose to change the static paradigm to an interactive one, develop systems that proactively ask questions to gather more information and respond reliably, and introduce an benchmark—MediQ—to evaluate question-asking ability in LLMs. MediQ simulates clinical interactions consisting of a Patient System and an adaptive Expert System; with potentially incomplete initial information, the Expert refrains from making diagnostic decisions when unconfident, and instead elicits missing details via follow-up questions. We provide a pipeline to convert single-turn medical benchmarks into an interactive format. Our results show that directly prompting state-of-the-art LLMs to ask questions degrades performance, indicating that adapting LLMs to proactive information-seeking settings is nontrivial. We experiment with abstention strategies to better estimate model confidence and decide when to ask questions, improving diagnostic accuracy by 22.3%; however, performance still lags compared to an (unrealistic in practice) upper bound with complete information upfront. Further analyses show improved interactive performance with filtering irrelevant contexts and reformatting conversations. Overall, we introduce a novel problem towards LLM reliability, an interactive MediQ benchmark and a novel question-asking system, and highlight directions to extend LLMs’ information-seeking abilities in critical domains.
@@ 1 Introduction
General-purpose large language models (LLMs) are designed to serve a broad audience by following instructions and providing the most likely and general answers (Brown et al., 2020; Ouyang et al., 2022). However, in high-stakes decision making scenarios such as clinical conversations, LLM assistants can be harmful if they provide general responses instead of gathering missing information to make informed decisions. As shown in Figure 1, standard medical question-answering (QA) tasks are formulated in a single-turn setup where all necessary information is provided upfront, and the model is not expected to interact with users (Jin et al., 2021; Pal et al., 2022; Jin et al., 2019; Hendrycks et al., 2020). This QA paradigm diverges from real-world scenarios, where users may provide incomplete information, and effective decision-making often requires an investigative process involving follow-up questions to clarify and gather necessary details (Trimble & Hamilton, 2016; Bornstein & Emler, 2001; Masic, 2022).
This gap between existing benchmarks and reality calls for a paradigm shift to designing systems adept at navigating high-stakes interactive scenarios. Focusing on clinical interactions where context is often incomplete, we introduce MediQ, an interactive benchmark for medical evaluation with dynamic information-seeking questions, to address limitations of static single-turn QA benchmarks (Figure 2). Unlike conventional systems, which assume that all necessary information is readily available, MediQ acknowledges the inherent uncertainty in medical consultations where a typical patient does not have the expertise to distill all necessary and relevant information they need to provide. To achieve this, MediQ comprises two components: a Patient system that simulates a patient and responds to follow-up questions, and an Expert system that serves as a doctor’s assistant and asks questions to the patient before making a medical decision. In this interactive clinical reasoning task, a successful information-seeking Expert should decide, at each turn, whether it has enough information to provide a confident answer; if not, it should ask a follow-up question.
@@
We convert two medical QA datasets, MedQA (Jin et al., 2021) and Craft-MD (Johri et al., 2023, 2024), into an interactive benchmark by parsing the patient records to only provide partial information in the beginning. We first develop and validate a Patient system that accurately answers Expert inquiries by retrieving the correct facts from the patient record. We then benchmark Expert systems based on state-of-the-art (SOTA) LLMs, including Llama-3 (Touvron et al., 2023), GPT-3.5 (Brown et al., 2020) and GPT-4 (OpenAI et al., 2024), to evaluate their proactive information seeking ability. It is striking that prompting these models to ask questions results in an 11.3% accuracy drop compared to starting with the same limited information and asking no questions, showing that adapting LLMs to interactive information-seeking settings is nontrivial. A key challenge is deciding when to ask a follow-up question instead of directly providing an answer. With confidence estimation strategies such as rationale generation and self-consistency, we improve Expert performance by 22.3%, although a 10.3% gap remains compared to an upper bound when full information is presented at once.
Our results show that while SOTA LLMs perform relatively well with complete information, they struggle to proactively seek missing information in a more realistic, interactive settings with incomplete initial information. By providing a modular, interactive benchmark, we hope to facilitate the development of reliable LLM assistants for complex decision-making in healthcare and other high-stakes domains. Our main contributions are:
@@
    We identify the critical problem of information-seeking questions in reliable interactive LLM assistants. We propose a paradigm shift and a practical conversion pipeline from standard single-turn benchmarks into interactive settings with incomplete initial information.
@@
    We develop the MediQ Benchmark to simulate more realistic clinical interactions between a Patient System and an Expert System. We rigorously develop and test the Patient System to benchmark any Expert’s information-seeking and clinical decision-making abilities.
@@
    We show that SOTA LLMs such as Llama-3-Instruct, GPT-3.5 and GPT-4 struggle at proactive information seeking, revealing a significant gap in this area.
@@
    We propose MediQ-Expert, our best Expert system with novel abstaining capabilities to reduce unconfident answers, to partially close the gap between the more realistic incomplete information setup and the existing full information setup.
Figure 2: The MediQ Benchmark. MediQ operationalizes a more realistic dynamic clinical interaction between a Patient system and an Expert system to evaluate info-seeking and question-asking.
@@ 2.2 The Expert System
Expert Task    The Expert system simulates the medical decision-making process of experienced clinicians, who seek additional patient information and iteratively update their differential diagnosis. The Expert system is first presented with a medical question and limited patient information. As each turn, it assesses whether the provided information is sufficient to answer the question. If the Expert system is unconfident, it can elicit evidence with a follow-up information-seeking question; otherwise, the Expert system deems the acquired information sufficient and provides a final answer. The performance of the Expert system is evaluated on the (1) efficiency of the conversation (number of follow-up questions) and (2) the accuracy of the final diagnosis.
@@ 5.2 How does abstention improve Expert system performance?
Figure 8: Effect of the abstention module on Expert system performance (efficiency & accuracy). (a) Accuracy over the number questions when the abstention response is not provided to the question generator context (NC) to isolate the effect of confidence thresholds on model performance. Each line is a different abstain method with increasing confidence thresholds at each point. Accuracy increases with # Qs as threshold increases. (b) Expert confidence estimation throughout the interaction with and without rationale generation (RG), averaged over self-consistency levels 1, 3, and 5. The expected calibration error (ECE) is shown in the legend. RG leads to more conservative and accurate confidence estimates. (c) Accuracy of Expert system variants with abstention response available (solid marker) vs. not available (hollow marker) to the question generator module. Including abstention response in question generator context improves accuracy, and is amplified with rationale generation (RG) and self-consistency (SC).
@@ 6 Related Work
Medical Question Answering Systems   Advancements in medical question answering (QA) systems have progressed from rule-based systems to LLM-powered agents. Notable medical QA benchmarks include MultiMedQA (Singhal et al., 2023a), which contains both multiple-choice and open-ended questions collected from various sources. To customize or adapt a general-purpose LLM to the medical domain, prior work often finetune the model on medical knowledge data such as PubMed (Bolton et al., 2022; Yasunaga et al., 2022; Wu et al., 2023a; Singhal et al., 2023a, b), or more recently on conversational medical datasets (Yunxiang et al., 2023; Han et al., 2023). Kim et al. (2024) further improves model performance on complex medical questions by dynamically forming multi-agent collaboration structures. Despite their proficiency in direct answer retrieval, the proactive information-seeking capability is not something these models are inherently designed to do. Our proposed methodological framework, MediQ, is designed to work as an overlay to these domain-specific models, enhancing them with the capability to actively seek additional information in a structured and clinically relevant manner.
Interactive Models and Agents   Interactive conversational models extend beyond the standard QA framework by engaging in a dialogue such as customer support and negotiation (Singh & Beniwal, 2022; Chakrabarti & Luger, 2015; He et al., 2018; Abdelnabi et al., 2023; JU et al., 2024), where iterative information gathering is crucial. Li et al. (2023) and Andukuri et al. (2024) attempt to use LLMs to elicit more information-rich human preference examples in everyday tasks. However, the application of these models in the medical domain remains limited (Li et al., 2021). Wu et al. (2023b) attempts to evaluate general-purpose LLMs and chain-of-thought reasoning on DDXPlus (Fansi Tchango et al., 2022), a rule-based synethtic patient dataset. Hu et al. (2024) navigates the information-seeking scenario as a search problem by developing a reward model guided by uncertainty and includes medical diagnosis as one of the tasks, but if limited to binary questions. Johri et al. (2023) observed a similar phenomenon where LLM-doctors cannot elicit complete patient information, but do not focus on improving the information-seeking ability. The system proposed by (Tu et al., 2024) performs a multitude of medical tasks but does not explore the crucial problem of abstention. Lastly, multi-agent and human-AI collaboration frameworks have shown impressive interactive performance (Zhou et al., 2024; Wu et al., 2024, 2023c; Deng et al., 2024; Lin et al., 2024), and can greatly benefit from our novel interactive abstention methods to seek additional information. Our work fills this gap via providing a benchmark, a dataset, and a framework to comprehensively studying information-seeking abilities in clinical decision-making, and most importantly, opens the door for future endeavors in this direction.
@@ 7 Conclusion
In this paper, we identify a significant gap in current LLMs’ capability to ask questions and proactively seek information in settings where personalization, precision, and reliability are critical. We propose a paradigm shift to interactive benchmarks by simulating more realistic clinical interactions where only partial information is provided initially by introducing MediQ. MediQ provides a benchmark to the community to evaluate the question-asking ability of LLMs, contributing towards developing reliable models. We showed that SOTA LLMs like Llama-3 and GPT-4 struggle to gather necessary information for accurate medical decisions. To address this problem, we presented MediQ-Expert—a novel Expert system with improved confidence judgment and medical expertise, substantially improving clinical reasoning performance. MediQ operationalizes interactive and explicit clinical reasoning processes, with added interpretability in the reasoning flow of language models and decision making. We encourage future research to extend MediQ to more diverse Patient systems, expand medical knowledge integration, and customize the interactions to better serve the healthcare community.
@@
One limitation is the scarcity of datasets that contain detailed patient information sufficient for a medical diagnosis which, to the best of our knowledge, was only met by MedQA and Craft-MD. The majority of available medical datasets are designed to test models’ medical domain knowledge. Second, the Patient system in our benchmark relies on a paid API; future work should establish an open-source Patient. Lastly, our evaluation framework, while designed to be more realistic, is still limited to the multiple-choice format. However, the flexibility of MediQ allows easy extensions into open-ended settings with appropriate datasets and well-defined conversation-level metric. Future work can focus on collecting a rich dataset in open-ended medical consultations and expanding the MediQ framework.
@@
If a similar framework is used in real-world applications, users and clinicians should be trained to prevent the over-reliance on technology that is liable to make mistakes, and to understand its privacy and fairness risks.
Acknowledgements
We gratefully acknowledge support from the University of Washington Population Health Initiative. SSL, SF, and YT are supported by the National Science Foundation under CAREER Grant No. IIS2142739, NSF grants No. IIS2125201, IIS2203097, and gift funding from Google, MSR, and OpenAI. PWK is supported by the Singapore National Research Foundation and the National AI Group in the Singapore Ministry of Digital Development and Innovation under the AI Visiting Professorship Programme (award number AIVP-2024-001). EP was supported by a Google Research Scholar award, Optum, NSF CAREER #2142419, a CIFAR Azrieli Global scholarship, a gift to the LinkedIn-Cornell Bowers CIS Strategic Partnership, and the Abby Joseph Cohen Faculty Fund.
@@ A.2.3 Patient System Variants: Fact-Select
"She says that, despite going to bed early at night, she is unable to fall asleep. She denies feeling anxious or having disturbing thoughts while in bed. Even when she manages to fall asleep, she wakes up early in the morning and is unable to fall back asleep. She says she has grown increasingly irritable and feels increasingly hopeless, and her concentration and interest
Atomic fact decomposition Prompt (cont’d) at work have diminished. The patient denies thoughts of suicide or death. Because of her diminished appetite, she has lost 4 kg (8.8 lb) in the last few weeks and has started drinking a glass of wine every night instead of eating dinner. She has no significant past medical history and is not on any medications."
@@ B.1 Initial Assessment
MediQ Expert – Initial Assessment Module Prompt (cont’d) Given the information from above, your task is to choose one of four options that best answers the inquiry.